OpenStudy (anonymous):

medal to check my answer- Which number make this equation true? v2 +10v +16=(v+8)(v+?) A.2 B.4 C.6 D.8 My answer B.4

jimthompson5910 (jim_thompson5910):

fill in the blanks 8 + ____ = 10 8 * _____ = 16

OpenStudy (anonymous):

Wow I was so off - its 2 can u help me with my last one ?

jimthompson5910 (jim_thompson5910):

it helps to check your answer by expanding it out

OpenStudy (anonymous):

I dont even know where to begin with this one.. The area of a garden is given by the trinomial g2 =2g-24 .The garden s width is g +4 What is the garden s width ? how does it break down? A g-2 b. g-6 c. g-8 d.g+2 I am getting g-6

jimthompson5910 (jim_thompson5910):

did you mean g^2 + 2g - 24 ?

jimthompson5910 (jim_thompson5910):

if so, then g^2 + 2g - 24 = (g+4)(g-6) which makes you correct

jimthompson5910 (jim_thompson5910):

actually no, i went too fast

jimthompson5910 (jim_thompson5910):

it should be -2g if that was the case

OpenStudy (anonymous):

ok how did u get -2g ?

jimthompson5910 (jim_thompson5910):

what is g2 =2g-24 supposed to be?

OpenStudy (anonymous):

It says the gardens lenght is g+4

jimthompson5910 (jim_thompson5910):

what is the original trinomial

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

what is the original trinomial?

OpenStudy (anonymous):

g2 -2 g -24 the garden lenght is g+4 Im getting( g+4)( g-6)

jimthompson5910 (jim_thompson5910):

good, g^2 - 2g - 24 = (g+4)(g-6)

OpenStudy (anonymous):

thank you again , i need more practice

jimthompson5910 (jim_thompson5910):

you're welcome